# LAVIB: A Large-scale Video Interpolation Benchmark

### v1

@@ 6 Conclusions
OOD over reduced train set sizes. Performance trends when removing the highest/lowest valued clips in OOD challenges given their metrics are reported in Table A9 and Table A9 for AFM and ALV. Similarly, Table A9 and Table A9 report results with training set reductions for ARMS and ARL. Across settings, the portions closer to the target domain; e.g. the top 30% for the low to high settings and bottom 30% for the high to low settings present the largest drop in performance when removed. In contrast, when portions of the data that are less similar to the target domain are removed the reductions in performance are marginal. This shows that VFI method trained on domain-specific videos cannot generalize as effectively.
@@
The introduced dataset primarily considers footage of landscapes, objects, nature, animals, and screen recordings. However, certain videos may include people. Scenes in which people appear are characterized by high camera motion, scene clutter, and partial visibility of faces briefly for a few seconds. Thus, it is believed that the risk of identification is low. In addition, the video segments from which the dataset is sourced are 1 second long, limiting the number of frames available. As videos are sourced from YouTube a list of the links to the original videos is also provided.

### v2

@@ 6 Conclusions and future directions
LAVIB further encourages exploring new avenues for efficiency improvements in future VFI works.
Frame-level quantization. A number of works have explored video inference acceleration through frame quantization for temporal redundancy reduction [1, 62]. Learning to truncate videos by varying quantization precision is important for the real-world applicability of methods in streams. LAVIB provides a diverse set of high-resolution videos with standardized frame rates that can be used both as a benchmark as well as a pre-training dataset.
Knowledge distillation. Transferring knowledge about the video structure can enable more efficient models. Several works [25, 35, 48] have distilled representations from teacher models trained on high-resolution videos. A natural extension of the proposed dataset would be its use for training high-resolution teacher models and evaluating low-resolution student models.
Salient frame sampling. The selection of informative frames has been another domain of interest for real-time video processing [66, 69, 70]. The standardized framerate of LAVIB can provide a robust benchmark for testing sampling approaches over different granularities.
Based on these adjacent video tasks, LAVIB can be imported and adapted as a general-purpose dataset and benchmark.
@@
OOD over reduced train set sizes. Performance trends when removing the highest/lowest valued clips in OOD challenges given their metrics are reported in Tab. A12 and Tab. A12 for AFM and ALV. Similarly, Tab. A12 and Tab. A12 report results with training set reductions for ARMS and ARL. Across settings, the portions closer to the target domain; e.g. the top 30% for the low to high settings and bottom 30% for the high to low settings present the largest drop in performance when removed. In contrast, when portions of the data that are less similar to the target domain are removed the reductions in performance are marginal. This shows that VFI method trained on domain-specific videos cannot generalize as effectively.
@@
Ethics and privacy. The introduced dataset primarily considers footage of landscapes, objects, nature, animals, and screen recordings. However, certain videos may include people. Scenes in which people appear are characterized by high camera motion, scene clutter, and partial visibility of faces that appear briefly for a few seconds. Thus, it is believed that the risk of identification is low. In addition, the video segments from which the dataset is sourced are 1 second long, limiting the number of frames available. As videos are sourced from YouTube a list of the links to the original videos is also provided.
Use. The dataset is distributed for open-source scientific projects under a Creative Common’s Attribution-NonCommercial-Share-Alike (CC BY-SA-NC 4.0). The dataset can be further shared, and adapted, but cannot be used for commercial applications. Adaptations or sharing of the dataset needs to be done under the same license†. ††footnotetext: †Clarifications on special use cases can be found in: https://creativecommons.org/licenses/by-nc-sa/4.0/deed.en